# The Converse Madelung Question Schrödinger Equation from Minimal Axioms

### v1

@@ 12.1 Information-geometric closure
endowing the $\psi$-representation with a Kähler structure. Under standard hypotheses on $V$ (e.g. $V$ Kato-small relative to $\Delta$ or $V\in L^{2}_{\mathrm{loc}}$ with form bounds), the Schrödinger operator is self-adjoint on $L^{2}$, so Stone’s theorem yields a one-parameter unitary group; see [19].
@@ 12.3 Variational self-consistency
$\partial_{t}(\rho v_{i})+\partial_{j}\Pi_{ij}\;=\;-\,\frac{\rho}{m}\,\partial_{i}V\;+\;\Bigl(\alpha-\frac{\hbar^{2}}{2m}\Bigr)\,\partial_{i}\!\left(\frac{\Delta\sqrt{\rho}}{\sqrt{\rho}}\right).$
Thus the local momentum balance closes if and only if $\alpha=\hbar^{2}/(2m)$. Any other coefficient leaves a nonzero residual divergence, so the Fisher scale is dynamically fixed by conservation.
@@ 12.4 Reversibility and the Doebner-Goldin class
$i\hbar\,\partial_{t}\psi=\left[-\frac{\hbar^{2}}{2m}\nabla^{2}+V\right]\psi+i\hbar D\left(\frac{\Delta\rho}{\rho}\right)\psi\qquad(\rho=|\psi|^{2}),$
@@
Placement. Within the DG family, the reversible sector is the $D=0$ corner singled out by Axiom V; this coincides with the Fisher-scaled Hamiltonian flow fixed by $\hbar^{2}=2m\alpha$.
@@
implies
$\frac{d}{dt}\,S_{\mathrm{Sh}}[\rho]=\frac{d}{dt}\int\rho\ln\rho\,dx=D\int\frac{|\nabla\rho|^{2}}{\rho}\,dx\;\geq\;0.$
@@
Any measured $\dot{S}_{\mathrm{Sh}}>0$ at fixed $V$ falsifies reversible dynamics; the reversible corner is $D=0$, coinciding with the Fisher-scaled Hamiltonian flow fixed by $\hbar^{2}=2m\alpha$.
Reversibility (Axiom V) holds only for $D=0$, which corresponds exactly to the Fisher coefficient fixed by $\hbar^{2}=2m\alpha$. Any nonzero diffusion or nonlinear gauge term generates irreversible or nonlinear evolution. The Fisher value is thus the reversible fixed point of the DG family.
12.5 Numerical and empirical falsifiers
@@ 13  Discussion and Implications
Within our admissible class, the Fisher-regularised Hamiltonian system is found as the only reversible completion of classical ensemble dynamics consistent with locality, conservation, Euclidean symmetry, and global $\mathrm{U}(1)$ phase symmetry. All other curvature forms break one or more of these constraints. Independent classifications of Hamiltonian structures are consistent with Fisher curvature being the unique first-derivative, positive scalar compatible with Euclidean covariance and reversibility within flat Dubrovin-Novikov brackets [15]. In this sense, within our axiomatic structural principles, quantum mechanics can be considered as a fixed point, with Planck’s constant $\hbar$ as a scale factor connecting the information-geometric curvature of probability space with the symplectic geometry of reversible dynamics.
@@ 13.1 Experimental falsification and universality tests
$Q[\rho]=\alpha\,\frac{\Delta\sqrt{\rho}}{\sqrt{\rho}},\qquad\alpha=\frac{\hbar^{2}}{2m},$
within the stated axioms, as the only correction compatible with both linearity and zero entropy production within our admissible class. Its absence or modification necessarily leads to dissipation or non-unitarity. This makes the result directly falsifiable.
If an analog physical system in condensed matter, optics, hydrodynamics, or emergent computation realises a reversible "quantum-like" dynamics within the stated axioms (local first-order, Hamiltonian, Euclidean-covariant, global $\mathrm{U}(1)$), then in its coarse-grained limit the effective Hamiltonian density should contain the Fisher term
@@
Strictly detached, and outside the paper’s axiomatic scope and claims; for diligence, future orientation and diagnostics only, we include a curvature guard (Test 10; App. E) - a fixed-background unit check of the covariant Fisher variation $\delta\mathcal{F}_{g}/\delta\rho=-\,\Box_{g}\sqrt{\rho}/\sqrt{\rho}$ (with the standard scalar freedom $\xi R\rho$) and a sketch showing how our residual protocol ports to a linear Klein-Gordon dispersion check when the discrete symbol is respected.
@@ Conclusion
Within our minimal axioms, the admissible reversible hydrodynamics on $(\rho,S)$ selects a single structure in the stated class. The bracket reduces to the canonical form, the only axiomatically compatible curvature is the Fisher functional, and the sole local gauge covariant complexification that linearises the flow is $\psi=\sqrt{\rho}\,e^{iS/\hbar}$ with $\hbar^{2}=2m\alpha$. In many body form, linearity with one complex structure is compatible with a single Planck constant through $\alpha_{i}=\hbar^{2}/(2m_{i})$ componentwise.
@@ Appendix A Appendix: Boundary Classes
$\delta\!\int_{\Omega}|\nabla R|^{2}dx=-\,2\!\int_{\Omega}\frac{\Delta R}{R}\,\delta\rho\,dx+\;2\!\oint_{\partial\Omega}(\nabla R\!\cdot\!n)\,\delta R\,\mathrm{d}\sigma.$
The boundary term vanishes under either $R|_{\partial\Omega}=0$ (Dirichlet) or $\nabla R\!\cdot\!n=0$ (Neumann), so $\delta\mathcal{F}/\delta\rho=-\Delta R/R$ is well defined in the weak sense on $\{\rho>0\}$.
@@ Appendix B Appendix: Counterexamples to Axioms
If $\{S,C\}\neq-1$, e.g. $\{S(x),\rho(y)\}=0$, the global phase symmetry fails and total probability $C=\!\int\rho\,dx$ is no longer conserved.
@@
Adding a preferred-direction term, e.g. $\beta\,\nabla S\!\cdot\!\hat{n}$, breaks isotropy and parity, violating energy covariance.
@@
Adding a diffusive term to continuity, $\partial_{t}\rho+\nabla\!\cdot(\rho\nabla S/m)=D\,\Delta\rho$ with $D>0$, yields $\frac{d}{dt}\int\rho\ln\rho\,dx=D\int|\nabla\rho|^{2}/\rho\,dx\geq 0,$ breaking time-reversal invariance. Equivalently, keeping a Hamiltonian form but modifying the bracket to $\{\rho,S\}=a_{0}(\rho)\delta$ violates Jacobi unless $a_{0}^{\prime}(\rho)=0$, so the flow exits the Poisson (reversible) class when $a_{0}$ varies with $\rho$.
@@
Any alternative $f(\rho)$ in $F[\rho]=\!\int f(\rho)|\nabla\rho|^{2}dx$ leaves a residual nonlinear term in the Hamilton-Jacobi equation that no local transformation can remove; only $f(\rho)\propto 1/\rho$ preserves reversibility.
@@ Appendix C Appendix: Jacobi Verification
Jacobi condition for local brackets. The special triple with two $S$’s is tautological once $\{S,S\}=0$: the identity $\{\rho,\{S,S\}\}+\mathrm{cyclic}=0$ reduces to $0\equiv 0$ and cannot constrain $a_{0}(\rho)$. For local, derivative-free brackets $\{u^{i},u^{j}\}=P^{ij}(u)\delta$, Jacobi reduces pointwise to
@@
With $P^{\rho S}(\rho)=a_{0}(\rho)$ the only nonzero derivative is $\partial_{\rho}a_{0}$, and the condition yields $a_{0}\,\partial_{\rho}a_{0}=0$, hence either $a_{0}\equiv 0$ or $a_{0}^{\prime}(\rho)=0$. Axiom II excludes $a_{0}\equiv 0$, so $a_{0}^{\prime}(\rho)=0$.
@@
Translation invariance removes explicit $x$-dependence, so $g_{ij}=g_{ij}(\rho)$. Rotation invariance for all profiles forces $g_{ij}(\rho)$ to transform as a scalar multiple of the identity (Schur’s lemma for the defining $SO(d)$ representation), hence $g_{ij}(\rho)=a(\rho)\delta_{ij}$. Global $\mathrm{U}(1)$ on $S$ and first-order locality exclude dependence on $S$ or derivatives of $\rho$. ∎
@@
Proof.
Translation invariance: for any $a\in\mathbb{R}^{d}$, invariance of $\mathcal{F}$ under $x\mapsto x+a$ implies $g_{ij}(\rho,x)$ cannot depend explicitly on $x$, so $g_{ij}=g_{ij}(\rho)$.
Rotation invariance: for any $R\in SO(d)$ and $\tilde{\rho}(x)=\rho(R^{-1}x)$, we require
$\int g_{ij}(\rho)\,\partial_{i}\rho\,\partial_{j}\rho\,dx=\int g_{ij}(\tilde{\rho})\,R_{ik}R_{j\ell}\,\partial_{k}\tilde{\rho}\,\partial_{\ell}\tilde{\rho}\,dx.$
Since the equality must hold for all $\rho$ and all $R$, Schur’s lemma for the defining representation of $SO(d)$ forces $g_{ij}(\rho)=a(\rho)\,\delta_{ij}$.
Phase generator decoupling: global $\mathrm{U}(1)$ acts as $S\mapsto S+\text{const}$ with $\rho$ fixed. Compatibility with the Hamiltonian structure and first order locality conflict with $a$ depending on derivatives of $\rho$ or on $S$. Thus $a=a(\rho)$ only.
Constancy: if $a$ varied with $\rho$, then under uniform rescalings on a connected set where $\nabla\rho\neq 0$ the tensor $a(\rho)\delta_{ij}$ would change value pointwise, contradicting strict rotational invariance of the quadratic form for arbitrary profiles. Equivalently, demanding that the quadratic form define the same inner product on gradient fields in every tangent space forces $a$ to be a constant. Hence $g_{ij}=\kappa\,\delta_{ij}$ with $\kappa>0$ by positivity. ∎
@@ Appendix D Appendix: Fisher-Curvature
The proportionality constant rescales the definition of $\alpha$ in Eq. (4.1), establishing Fisher necessity. A numerical verification of this Euler-Lagrange identity is provided in Test 5 (code archive, Appendix E).
@@ 3. Test 3:
    Supports result: $D=0$ reversible; $D>0$ gives $\dot{H}=D\,I_{F}$ and matches PDE to $10^{-14}$. Script: 3_entropy_production_DG.py. Notes: Verifies $dS_{\mathrm{Sh}}/dt=D\!\int|\nabla\rho|^{2}/\rho\,dx\geq 0$ and $\dot{H}$ identity.
@@ 8. Test 8:
    Supports result: Unique local complexifier $\psi=\sqrt{\rho}\,e^{iS/\hbar}$ with $\alpha=\hbar^{2}/(2m)$. Script: 8_complexifier_rigidity.py. Notes: Scans local $F(\rho),G(S,\rho)$ ansätze; only polar map linearises with $\alpha=\kappa^{2}/(2m)$.
@@ 10. Test 10:
    Records on fixed backgrounds the geometric identity $\delta\mathcal{F}_{g}/\delta\rho=-\,\Box_{g}\sqrt{\rho}/\sqrt{\rho}$ and the standard scalar freedom $\xi R\rho$; shows that our residual methodology transports to linear Klein-Gordon when the discrete symbol is respected. No curved dynamics are claimed. As noted strictly earlier, this test regards Scope and limitations only. It makes no claims regarding the central thesis.
@@ Appendix F Appendix: Conservation of Probability
For $\rho\!\geq 0$ with $\sqrt{\rho}\in H^{1}(\Omega)$, $\mathcal{F}[\rho]=\int|\nabla\sqrt{\rho}|^{2}dx$ is finite and strictly convex; its Euler-Lagrange derivative is well defined in $H^{-1}_{\mathrm{loc}}$. Any other $f(\rho)$ produces mixed $|\nabla\rho|^{2}$ terms in the Hamilton-Jacobi equation and destroys reversibility.
Direct evaluation for Gaussian $\rho(x)=e^{-x^{2}/\sigma^{2}}$ shows that $\big\|\delta\mathcal{F}/\delta\rho+4C\,\Delta\sqrt{\rho}/\sqrt{\rho}\big\|_{L^{2}}$ vanishes to machine precision only for $f=C/\rho$, indicating the analytic condition. Additionally verified by code in Appendix E.

### v2

@@ 12.1 Information-geometric closure
endowing the $\psi$-representation with a Kähler structure. Under the function-space and Kato-smallness hypotheses of Sec. 9, the Schrödinger operator is self-adjoint on $L^{2}$, so Stone’s theorem yields a one-parameter unitary group; see [19].
@@ 12.3 Variational self-consistency
$\partial_{t}(\rho v_{i})+\partial_{j}\Pi_{ij}\;=\;-\,\frac{\rho}{m}\,\partial_{i}V\;+\;\frac{\rho}{m}\Bigl(\alpha-\frac{\hbar^{2}}{2m}\Bigr)\,\partial_{i}\!\left(\frac{\Delta\sqrt{\rho}}{\sqrt{\rho}}\right).$
Thus the local momentum balance closes if and only if $\alpha=\hbar^{2}/(2m)$. Any other coefficient leaves a nonzero residual force density, so the Fisher scale is dynamically fixed by conservation.
@@ 12.4 Reversibility and the Doebner-Goldin class
$i\hbar\,\partial_{t}\psi=\left[-\frac{\hbar^{2}}{2m}\nabla^{2}+V\right]\psi+i\frac{\hbar D}{2}\left(\frac{\Delta\rho}{\rho}\right)\psi\qquad(\rho=|\psi|^{2}),$
@@
Placement. Within the DG family, the reversible sector is the $D=0$ corner singled out by Axiom V; this coincides with the Fisher-scaled Hamiltonian flow with Fisher coefficient $\alpha=\hbar^{2}/(2m)$.
The diffusive branch of this family thus acts as a controlled way to leave the reversible corner by relaxing the reversibility axiom, and nonlinear gauge-related forms remain visible in our framework through their failure of the superposition and entropy diagnostics introduced later.
@@
implies that the Shannon entropy
$S_{\mathrm{Sh}}[\rho]:=-\int\rho\ln\rho\,dx$
satisfies
$\frac{d}{dt}\,S_{\mathrm{Sh}}[\rho]=D\int\frac{|\nabla\rho|^{2}}{\rho}\,dx\;\geq\;0.$
@@
Any measured $\dot{S}_{\mathrm{Sh}}>0$ at fixed $V$ falsifies reversible dynamics. Within the DG family, reversibility (Axiom V) holds if and only if $D=0$, and this $D=0$ corner coincides with the Fisher-scaled Hamiltonian flow with $\alpha=\hbar^{2}/(2m)$. Any nonzero diffusion or nonlinear gauge term generates irreversible or nonlinear evolution, so the Fisher value is the reversible fixed point of the DG family.
12.5 Numerical and empirical falsifiers within class
@@ 13  Discussion and Implications
Within our admissible class, the Fisher-regularised Hamiltonian system is found as the only reversible completion of classical ensemble dynamics consistent with locality, conservation, Euclidean symmetry, and global $\mathrm{U}(1)$ phase symmetry. All other curvature forms break one or more of these constraints. Independent classifications of Hamiltonian structures are consistent with Fisher curvature being the unique first-derivative, positive scalar compatible with Euclidean covariance and reversibility within flat Dubrovin-Novikov brackets [15].
In this sense, within our axiomatic structural principles, quantum mechanics can be considered as a fixed point, with Planck’s constant $\hbar$ as a scale factor connecting the information-geometric curvature of probability space with the symplectic geometry of reversible dynamics.
@@
In this precise sense the construction does not merely display one possible realisation of quantum dynamics, but exhausts the reversible options available under the stated structural principles.
@@ 13.1 Experimental falsification and universality tests
$Q[\rho]=-\,\alpha\,\frac{\Delta\sqrt{\rho}}{\sqrt{\rho}},\qquad\alpha=\frac{\hbar^{2}}{2m},$
within the stated axioms, as the only correction compatible with both linearity and zero entropy production within our admissible class.
Its absence or modification necessarily leads to dissipation or non-unitarity. This makes the result directly falsifiable.
If an analog physical system in condensed matter, optics, hydrodynamics, or emergent computation realises a reversible quantum-like dynamics within the stated axioms (local first-order, Hamiltonian, Euclidean-covariant, global $\mathrm{U}(1)$), then in its coarse-grained limit the effective Hamiltonian density should contain the Fisher term
@@
This final test is strictly detached from, and lies outside, the paper’s axiomatic scope and claims. For diligence, future orientation, and diagnostics only, we include a curvature guard (Test 10; App. E), a fixed-background unit check of the covariant Fisher variation $\delta\mathcal{F}_{g}/\delta\rho=-\,\Box_{g}\sqrt{\rho}/\sqrt{\rho}$ (with the standard scalar freedom $\xi R\rho$) and a sketch showing how our residual protocol ports to a linear Klein-Gordon dispersion check when the discrete symbol is respected.
@@ Conclusion
Within our minimal axioms, the admissible reversible hydrodynamics on $(\rho,S)$ selects a single structure in the stated class. We do not assert uniqueness beyond this class, only that within it the Fisher-Schrödinger structure is forced. The bracket reduces to the canonical form, the only axiomatically compatible curvature is the Fisher functional, and the sole local gauge covariant complexification that linearises the flow is $\psi=\sqrt{\rho}\,e^{iS/\hbar}$ with $\hbar^{2}=2m\alpha$. In many-body form, linearity with one complex structure is compatible with a single Planck constant through $\alpha_{i}=\hbar^{2}/(2m_{i})$ componentwise.
@@
We have not addressed Born’s rule, measurement, or quantum field theory, and we have not attempted to derive the empirical linearity of quantum mechanics from a more microscopic theory. Our result is conditional: given a DN local reversible hydrodynamics that admits a projectively linear complex representation, the admissible theory space collapses to the Fisher scaled Schrödinger equation. In this sense the usual linear dynamics are classified rather than simply postulated.
@@ Appendix A Appendix: Boundary Classes
Neumann boundaries model reflecting walls and in general break global Galilean boost symmetry; accordingly, we do not use Neumann data in the Bargmann algebra check of Appendix J.
@@
$\delta\!\int_{\Omega}|\nabla R|^{2}dx=-\!\int_{\Omega}\frac{\Delta R}{R}\,\delta\rho\,dx+\;2\!\oint_{\partial\Omega}(\nabla R\!\cdot\!n)\,\delta R\,\mathrm{d}\sigma.$
The boundary term vanishes under either $R|_{\partial\Omega}=0$ (Dirichlet) or $\nabla R\!\cdot\!n=0$ (Neumann), so $\delta\mathcal{F}/\delta\rho=-\Delta R/R$ is well defined in the weak sense on $\{\rho>0\}$. Here $\mathcal{F}[\rho]=\int_{\Omega}|\nabla\sqrt{\rho}|^{2}dx$ is the Fisher regulariser in the $R$-variable; in the main text the regularising Hamiltonian always appears in the equivalent form $\mathcal{H}_{\mathrm{reg}}=\int\alpha\,|\nabla\sqrt{\rho}|^{2}dx$.
@@ Appendix B Appendix: Counterexamples to Axioms
If $\{S,C\}\neq-1$, e.g. $\{S(x),\rho(y)\}=0$, the global phase symmetry fails and the special role of $C=\!\int\rho\,dx$ as phase generator decouples from probability flow; for generic Hamiltonians the total probability $C$ will not be conserved.
@@
Adding a preferred-direction term, e.g. $\beta\,\nabla S\!\cdot\!\hat{n}$, breaks isotropy and parity and hence violates Euclidean covariance of the energy.
@@
Adding a diffusive term to continuity, $\partial_{t}\rho+\nabla\!\cdot(\rho\nabla S/m)=D\,\Delta\rho$ with $D>0$, yields
$\frac{d}{dt}\Bigl(-\!\int\rho\ln\rho\,dx\Bigr)=D\int|\nabla\rho|^{2}/\rho\,dx\geq 0,$
so the Shannon entropy $S_{\mathrm{Sh}}[\rho]:=-\int\rho\ln\rho\,dx$ increases monotonically and time-reversal invariance is broken. Equivalently, keeping a Hamiltonian form but modifying the bracket to $\{\rho,S\}=a_{0}(\rho)\delta$ changes the role of $\mathcal{C}=\!\int\rho\,dx$ as a phase generator; any non-constant $a_{0}(\rho)$ moves the system outside the canonical Hamiltonian class used here.
@@
We restrict attention to convex local regularisers of the form $F[\rho]=\int f(\rho)\,|\nabla\rho|^{2}\,dx$ with $f(\rho)>0$. Within this class we will show, via Proposition 2.3 (proved in Appendix D), that any alternative $f(\rho)$ leaves a residual nonlinear term in the Hamilton-Jacobi equation that no local transformation can remove, and that the unique choice compatible with exact projective linearity after a local complexifier is therefore $f(\rho)\propto 1/\rho$; no Fisher form is presupposed at the axiomatic level.
@@ Appendix C Appendix: Jacobi Verification
Jacobi condition for local brackets. The special triple with two $S$’s is tautological once $\{S,S\}=0$: the identity $\{\rho,\{S,S\}\}+\mathrm{cyclic}=0$ reduces to $0\equiv 0$ and cannot constrain $a_{0}(\rho)$. For local, derivative-free brackets $\{u^{i},u^{j}\}=P^{ij}(u)\delta$ the Jacobi identity can be written as the vanishing of the Schouten bracket
@@
In the present two-component setting the only potentially nonzero component is $P^{\rho S}(\rho)=a_{0}(\rho)=-P^{S\rho}$, and for any smooth $a_{0}(\rho)$ the Schouten bracket vanishes identically. Thus Jacobi places no further restriction on $a_{0}$ beyond those already implied by Axiom II, which then fixes $a_{0}$ to be a nonzero constant that can be absorbed into a rescaling of $S$, in agreement with Lemma 3.3 in the main text.
@@
Translation invariance removes explicit $x$-dependence, so $g_{ij}=g_{ij}(\rho)$. Rotation invariance for all profiles forces $g_{ij}(\rho)$ to transform as a scalar multiple of the identity (Schur’s lemma for the defining $SO(d)$ representation), hence $g_{ij}(\rho)=a(\rho)\delta_{ij}$. Global $\mathrm{U}(1)$ on $S$ and first-order locality exclude any dependence of $g_{ij}$ on $S$ or on derivatives of $\rho$.
@@ Appendix D Appendix: Fisher-Curvature
Comparing with the Fisher potential $Q_{\alpha}(\rho)=-\alpha\,\Delta\sqrt{\rho}/\sqrt{\rho}$ shows that $\alpha=4C$. Writing $C=\alpha/4$ we can equivalently parameterise the functional as
$F[\rho]=\int\frac{\alpha}{4}\,\frac{|\nabla\rho|^{2}}{\rho}\,dx=\alpha\int|\nabla\sqrt{\rho}|^{2}dx,$
which is the normalisation used in the main text. In particular, the complexifier analysis in Appendix K fixes $\alpha=\kappa^{2}/(2m)$, and identifying $\kappa=\hbar$ yields the Schrödinger value $\alpha=\hbar^{2}/(2m)$.
A numerical verification of this Euler-Lagrange identity is provided in Appendix E.
@@ 3. Test 3:
    Supports result: $D=0$ reversible; $D>0$ gives $\dot{H}=D\,I_{F}$ and matches the DG PDE to $10^{-14}$. Script: 3_entropy_production_DG.py. Notes: Verifies $dS_{\mathrm{Sh}}/dt=D\!\int|\nabla\rho|^{2}/\rho\,dx\geq 0$ for $S_{\mathrm{Sh}}[\rho]:=-\int\rho\ln\rho\,dx$ and confirms the energy identity $\dot{H}=D\,I_{F}$ for the Fisher information $I_{F}=\int|\nabla\rho|^{2}/\rho\,dx$.
@@ 8. Test 8:
    Supports result: Unique local complexifier $\psi=\sqrt{\rho}\,e^{iS/\hbar}$ with $\alpha=\hbar^{2}/(2m)$. Script: 8_complexifier_rigidity.py. Notes: Scans local $F(\rho),G(S,\rho)$ ansätze; only the polar map linearises the reversible flow, with Fisher coefficient $\alpha=\kappa^{2}/(2m)$ and $\kappa$ numerically recovered as $\kappa=\hbar$.
@@ 10. Test 10:
    Records on fixed backgrounds the geometric identity $\delta\mathcal{F}_{g}/\delta\rho=-\,\Box_{g}\sqrt{\rho}/\sqrt{\rho}$ and the standard scalar freedom $\xi R\rho$; shows that our residual methodology transports to the linear Klein-Gordon setting when the discrete symbol is respected. No curved dynamics are claimed. As noted earlier, this test concerns scope and limitations only and makes no claims regarding the central thesis.
@@ Appendix F Appendix: Conservation of Probability
For $\rho\!\geq 0$ with $\sqrt{\rho}\in H^{1}(\Omega)$, $\mathcal{F}[\rho]=\int|\nabla\sqrt{\rho}|^{2}dx$ is finite and strictly convex; its Euler-Lagrange derivative is well defined in $H^{-1}_{\mathrm{loc}}$. Any other $f(\rho)$ produces mixed $|\nabla\rho|^{2}$ terms in the Hamilton-Jacobi equation and obstructs the linear reversible completion within our admissible class.
Direct evaluation for Gaussian $\rho(x)=e^{-x^{2}/\sigma^{2}}$ shows that $\big\|\delta\mathcal{F}/\delta\rho+4C\,\Delta\sqrt{\rho}/\sqrt{\rho}\big\|_{L^{2}}$ vanishes to machine precision only for $f=C/\rho$, numerically confirming the analytic condition derived in Appendix D. See also the implementation in Appendix E.